OpenStudy (anonymous):

Find the volume of a cube that has a total area of 96 sq. in. V = cu. in. A prism of height 12" has a rhombus with diagonals 6" and 8" for a base. Find the volume and the lateral area. V = cu. in. L.A. = sq. in.

OpenStudy (anonymous):

find the area of one face of the cube = 96/6 = 16 sq in now u can find length of one side volume = this length cubed

OpenStudy (anonymous):

volume of pyramid = 1/3 * area of the base * height

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

each face of a cube is a square whose area is 16 sq in - so how do you find length of one side? the area = length of side squared = 16: - in math form L * L = 16 so whats the value of L?

OpenStudy (anonymous):

what number when multiplied by itself gives 16?

OpenStudy (anonymous):

u still with us?

OpenStudy (anonymous):

hope i was'nt too forceful !!!!